OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - 2y - z, if x = 2, y = 3, and z = -4

OpenStudy (anonymous):

just plug the values of x y and z and get a number

OpenStudy (anonymous):

4(2) -2(3) -(-4) = 8 - 6 + 4 = 6